The recommended childrens dosage of vibramycin is 2mg/LBS of body weight divided in 2 doses on first day. The child weighs 40 LBS. Vibramycin syrup is available in 50mg/5mL. Amount for 1 dose is:
Answer #1

The total dosage for the first day is 2 mg \times 40 lbs = 80 mg is divided into two doses

Each dose is 2 \div 80 = 40 mg for single divided dose

The required medication = required dose \times available volume \div available dose

                                        = 40 \times 5 \div 50

                                        = 4

The amount of medication for a single dose in two divided doses for the first day is 4 ml each dose.